OpenStudy (richgall____):

LAST ONE A triangle has sides 8 centimeters, 10 centimeters, and 15 centimeters. Is the triangle a right triangle?

OpenStudy (richgall____):

DONT Two sides have to be equal

OpenStudy (retireed):

No, because 3 cm, 4 cm and 5 cm is a right triangle

OpenStudy (richgall____):

so is this one a right triangle?

OpenStudy (retireed):

Use this equation the Pythagorean theorem: \[a ^{2} + b ^{2} = c ^{2}\] to prove or disprove if the triangle is a right triangle.

OpenStudy (richgall____):

so I do i plug in all 3 numbers ?

OpenStudy (richgall____):

@pooja195

OpenStudy (retireed):

In a right triangle, the longest side is the hypotenuse which is the angle opposite the right angle. So plug the shorted two sides into the a squared and b squared side and see if they equal the longer side squared. As in the 3, 4 and side triangle 3^2 + 4^2 = 5^2 9 + 16 = 25 is a right triangle

OpenStudy (richgall____):

ok

OpenStudy (richgall____):

8^2=64 10^2=100 15^=225